OpenStudy (anonymous):

2a + 2b = c for b How do I solve for the indicated variable?

jimthompson5910 (jim_thompson5910):

\[\large 2a + 2b = c\] \[\large 2b = c-2a\] \[\large b = \frac{c-2a}{2}\]